OpenStudy (anonymous):

The present cost of a car is $9,000. The cost of the car depreciates at the rate of $1000 per year

OpenStudy (anonymous):

Write a function to show the cost of the car f(t) after t years i put :f(x) = 1000(x) - 9000

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Not quite...see the constant would be the cost of the car...and that would be 9000 ....that would be added to the function... the value DEPRECIATES at a rate of 1000 per year...that would be our variable and it would be negative. f(x) = -1000x + 9000

OpenStudy (johnweldon1993):

The question then states to show what the function would be after 't' years... "that means replacing our 'x' with 't' ... so f(t) = -1000t + 9000

OpenStudy (anonymous):

Part B: What is the total cost of the car after 3 years? (4 points) Part C: If the cost of the car was $500 less than the present cost, what would be the cost of the car after 5 years? (2 points)

OpenStudy (johnweldon1993):

Alright so part B would just be replacing 't' with (3) so f(3) = -1000(3) + 9000 f(3) = -3000 + 9000 f(3) = 6000 <---that would be the cost of the car after 3 years

OpenStudy (johnweldon1993):

If the car was 500 less than the present (so 9000 - 500 = 8500) what would the cost be after 5 years? so we have a new function f(t) = -1000t + 8500 replace 't' with 5 f(5) = -1000(5) + 8500 f(5) = -5000 + 8500 f(5) = 3500 That would be our cost after 5 years if the original price was 8500 :)